OpenStudy (anonymous):

Factor to find all possible values for x : 2x2^ + 9x + 4 =0

OpenStudy (jhannybean):

Is it \(\large 2x^2 +9x +4 =0\)?

OpenStudy (anonymous):

yes @Jhannybean

OpenStudy (anonymous):

do you know the quadratic formula? if not (and even if you do), you may want to check this out... http://www.youtube.com/watch?v=z6hCu0EPs-o

OpenStudy (anonymous):

\[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

I honestly think I know the answer which is x = -5/2 & x = -1 @pgpilot326

OpenStudy (anonymous):

i get something different...

OpenStudy (anonymous):

O.O how so?

OpenStudy (dan815):